# Tractable Gaussian Phase Retrieval with Heavy Tails and Adversarial Corruption with Near-Linear Sample Complexity

### v1

@@ 2 Techniques
Prior techniques for phase retrieval with strong adversarial corruption and heavy-tailed noise (1.2): Naturally, the main algorithmic principle of [7] and [23] is to use a robust spectral initialization followed by a robust gradient descent procedure. We first describe their robust gradient descent procedure: they express the gradient as a mean of a random variable whose samples can be obtained from the 1.2, and leverage recent algorithmic breakthroughs for robust mean estimation; this idea has appeared in several recent works [53, 41, 36]. For robust spectral initialization, [7] have the insight that the top eigenvector of $\operatorname{Cov}(ya)$ also suffices for spectral initialization and use the exponential time robust covariance estimator [51] to estimate $\operatorname{Cov}(ya)$, and find the top eigenvector of the estimate. On the other hand [23] (who focuses on the $\sigma=0$ case) showed that given sufficient samples, robustly finding the top eigenvector of $\mathbb{E}[yaa^{T}]$ can be reduced to a convex Ky-fan-2 norm optimization problem, which can be converted into a packing semidefinite program for which nearly-linear time solvers exist.  [^6] Their reduction for near-linear sample complexity appears to have a bug; however, their reduction works if the sample complexity is allowed to be more than $\tilde{O}(n)$ (but still a polynomial in $n$), we elaborate on this in Section 4.2.
@@ 2.1 Our technical contribution:
Comparison of our technique with prior use of truncation+filtering. While truncation and filtering are classical tools in robust statistics, their successful interplay is highly problem-dependent. Several examples illustrate this point: for instance, [19]’s robust sparse mean estimation proceeds by projecting (i.e. truncating) means to an $\ell_{\infty}$ ball around the median-of-means estimator, followed by the filtering procedure of [2]. Catoni and Olivier [11] use a customized truncation-based estimator for heavy-tailed mean estimation. The work [56] addresses the robust sparse regression problem by first using coordinate-wise truncation together with a filtering step to remove adversarially corrupted design vectors, and then they perform regression by minimizing the Huber loss, which is essentially a truncated $\ell_{2}$ loss, to handle corruption in the responses. On the other hand, the proof attempt in the (incorrect) submission [23] uses truncation together with concentration inequalities, which is conceptually similar to the truncation-plus-filtering framework; from this viewpoint, the issues in their argument can be interpreted as a failed attempt to apply a truncation-based method. These examples highlight that the correct use of truncation and filtering is subtle and depends critically on the specifics of the problem: over-truncation can destroy essential structure, while under-truncation can render filtering ineffective. We truncate the vectors $ya$ using a norm-based cutoff determined by the input parameters, chosen large enough to keep the top eigenvector of the covariance of the truncated $ay$ close to the top eigenvector of the true covariance, yet small enough to apply the robust PCA method of [38]. Furthermore, unlike earlier truncation-based phase retrieval methods—which typically truncate only on the responses—our method truncates both the responses and the sensing vectors. We also show that our method remains valid even when the noise has a nonzero mean, which requires robustly solving a special case of blind deconvolution. Our work provides a principled and effective initialization scheme using truncation + filtering that may naturally extend to other robust non-convex estimation settings.
@@ 3.2 Robust principal component analysis
Note that Algorithm 2 from [38] is actually designed to return the robust top eigenvector in the case $k=1$ but we can estimate the corresponding eigenvalue also. Look at the first paragraph in Section D.1 for more details. We provide a rigorous proof for Equation 4 in Section D.1.
@@ 4 Robust Phase Retrieval with zero mean noise
$\frac{dist(x_{P},x^{*})}{\|x^{*}\|}$ $\leq C_{5}\exp\!\left(-C_{6}P(1-\sqrt{\epsilon})\right)$
$\quad+C_{5}\frac{\sigma}{\|x^{*}\|^{2}}\left(\sqrt{\tfrac{n\log n}{\tilde{m}}}+\sqrt{\tfrac{\log(1/\delta)}{\tilde{m}}}+\sqrt{\epsilon}\right)$ (5)
@@
$\frac{dist(x_{P},x^{*})}{\left\|x^{*}\right\|}\leq C_{3}\exp\left(-C_{4}P(1-\sqrt{\epsilon})\right)$
$+C_{3}\frac{\sigma}{\|x^{*}\|^{2}}\left(\sqrt{\frac{n\log n}{\tilde{m}}}+\sqrt{\frac{\log(1/\delta)}{\tilde{m}}}\right)+C_{3}\frac{\sigma}{\|x^{*}\|^{2}}\sqrt{\epsilon}.$
@@ 4.1 Robust Spectral initialization
(Robust spectral initialization for phase retrieval.) Under the setting and notation of Algorithm 1, if $\epsilon\leq C_{1}(r_{up}^{2}+1)^{-2}$, $m_{1}\geq\log(2/\delta)(1+r_{up}^{2})$ and $m_{2}\geq C_{3}\left(n(r_{up}^{2}+1)\log(2n/(\delta\epsilon))/\epsilon\right)$. Then, with probability at least $1-\delta$, $dist\left(x_{0},x^{*}\right)\leq\left\|x^{*}\right\|/9.$ The time complexity of Algorithm 1 is $O(T_{\mathrm{rob\text{-}mean}}(m_{1},1)+m_{2}^{2}n).$
@@ 4.2 Connection between stability-based robust PCA and robust phase retrieval
By Proposition 4.7, we conclude that to obtain nearly-linear time algorithms with sample complexity $\tilde{O}(n)$ for robust phase retrieval, it suffices to prove the stability condition (Definition 4.6) for the empirical distribution with $\widetilde{O}(n)$ truncated samples, specifically, to answer in the affirmative the following question: 777[23] claims a proof of the open question for the noiseless case ($\zeta=0$) in their manuscript. However, the validity of their argument appears to rely critically on the last inequality on page 14, which claims that $\operatorname{Pr}_{x\sim\mathcal{N}(0,1)}(x^{4}\geq\ln z/c)\leq z^{-3}$ holds for all $z>1$ for some constant c; however, this is false, and more generally the integral $\int^{\infty}_{\tilde{\tau}}\operatorname{Pr}_{x\sim\mathcal{N}(0,1)}(x^{4}\geq\ln z/c)dz$ diverges for all $\tilde{\tau}$ and constants $c>0$.
@@
[…]
@@ D.1 Proof of Equation 4
As the Remark 3.3 suggests that Algorithm 2 [38] has been designed to retrieve the top eigenvector for the case $k=1$. Specifically, Algorithm 2 takes as input the corrupted set $T=\{X_{j}X_{j}^{\top}\}_{i=1}^{m}$, together with additional parameters such as the corruption level and failure probability. It then applies a filtering procedure (Algorithm 3, Double Filtering, from [38]) to remove points in $T$ that are likely to be adversarially corrupted. The average of the remaining points in $T$ is computed, which we denote by $\hat{M}$. Finally, the top eigenvector of $\hat{M}$ is returned as an estimate of the robust PCA solution, and the guarantee is given by Equation 3. In the last step, one may also compute the corresponding eigenvalue of this top eigenvector, which serves as an estimate of the leading eigenvalue of $\Sigma_{r}$. While Proposition 2.6 from [38] provides guarantees only for the eigenvector estimate (Equation 3) returned by Algorithm 2, the corresponding eigenvalue estimate can also be analyzed using the results in [38]. Its guarantee is stated in Equation 4. Here , we rigorously proof Equation 4. Before proving Equation 4, we state a lemma from [38] that will be used in the proof. The lemma is as follows:
@@ Appendix E Robust Phase retrieval under Zero-Mean Noise
$\mathbb{E}\left[\left(a^{\top}x^{*}\right)^{8}\left(v^{\top}a\right)^{4}\right]\leq 10395\,\|x^{*}\|^{4}(v^{\top}x^{*})^{4}+\|x^{*}\|^{8}(315+5670)\,\|v\|^{4}\leq 16380\,\|x^{*}\|^{8}\|v\|^{4}=O(\|x^{*}\|^{8}\|v\|^{4}),$ (18)
@@
$\|\mathbb{E}[XX^{\top}\mathbb{I}_{\|X\|\geq\tau}]\|_{\mathrm{op}}$ $=\sup_{\|v\|=1}\mathbb{E}[\langle v,X\rangle^{2}\mathbb{I}_{\|X\|\geq\tau}]\overset{(a)}{\leq}\sup_{\|v\|=1}\sqrt{p}\sqrt{\mathbb{E}[\langle v,X\rangle^{4}]}\leq\sup_{\|v\|=1}\sqrt{p}\sqrt{\mathbb{E}[8(\zeta^{4}+\langle a,x^{*}\rangle^{8})\langle a,v\rangle^{4}]}$
$\overset{(b)}{\leq}\sqrt{p}\sqrt{8(3K_{4}^{4}+16380\|x^{*}\|^{8})}=O\left(\sqrt{p}\sqrt{(K_{4}^{4}+\|x^{*}\|^{8})}\right)=O(\sqrt{p}(K_{4}^{2}+\|x^{*}\|^{4})),$
@@
$\|\hat{u}-u\|^{2}=O\left(\epsilon(r_{up}^{2}+1)+(r_{up}^{2}+1)\sqrt{\epsilon}\right),$
@@
$=(\lambda_{1}-\lambda_{2})/2\|\hat{u}-u\|^{2}\geq 5\|x^{*}\|^{4}\|\hat{u}-u\|^{2}.$
@@
$\|\hat{u}-u\|^{2}=O\left(\epsilon\left(1+\frac{\sigma^{2}}{\|x^{*}\|^{4}}\right)+73(r_{up}^{2}+1)\sqrt{\epsilon}\right)=O\left(\epsilon(r_{up}^{2}+1)+(r_{up}^{2}+1)\sqrt{\epsilon}\right).$
@@
$\|\hat{u}-u\|^{2}\leq O\left(\epsilon\left(1+\frac{\sigma^{2}}{\|x^{*}\|^{4}}\right)+73(r_{up}^{2}+1)\sqrt{\epsilon}\right)=O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)\leq 1/27$
@@
$dist(\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{12\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27}.$
@@
$dist\left(x_{0},x^{*}\right)$ $=dist\left(\hat{\|x^{*}\|}\hat{u},x^{*}\right)$
$\leq\|\hat{\|x^{*}\|}\hat{u}-\|x^{*}\|\hat{u}\|+\|\|x^{*}\|\hat{u}-\|x^{*}\|u\|+dist\left(\|x^{*}\|u,x^{*}\right)$
$=|\hat{\|x^{*}\|}-\|x^{*}\||+\|x^{*}\|\|\hat{u}-u\|+dist\left(\|x^{*}\|u,x^{*}\right)$

### v2

@@ 1 Introduction
Motivation for heavy-tailed noise and adversarial corruption: Real-world data are rarely clean and typically contain multiple sources of noise. Broadly, these noises can be classified into two categories: uncorrelated noise across samples and correlated noise across the dataset. In the worst case, uncorrelated noise can be modeled as stochastic heavy-tailed noise, while correlated noise is often modeled through the lens of a strong adversary (1.1) that can inspect the entire dataset and arbitrarily corrupt a constant fraction of samples. Such partial corruption models are well motivated by practical applications. For instance, wireless sensor networks [64, 55, 62, 12] and IoT systems consist of many spatially distributed, low-power devices that continuously sense, collect, and transmit environmental data. These multi-sensor platforms are widely deployed in settings such as smart grids, industrial IoT, and seismic monitoring arrays, where their measurements are aggregated to form global statistics for monitoring, inference, and decision-making. In such large-scale deployments, typically only a small subset of sensors is compromised [64, 68], due to factors such as limited physical access, partial network penetration, or firmware vulnerabilities that affect only certain devices. The majority of sensors remain honest. Consequently, the data arriving at the central aggregator naturally fit a model in which a fraction of the measurements are adversarial, while the rest are clean, resulting in a constant fraction adversarial contamination model (1.1). This setting directly motivates the framework of robust statistics, which has been extensively studied; see, for example, an entire book [22] is devoted to this contamination model, and numerous fundamental problems have been studied under it, including robust mean estimation [40], covariance estimation [42, 43, 27], PCA [41, 24], linear regression [14], and non-convex optimization [7, 44]. Importantly, phase retrieval itself is a highly non-convex inverse problem in which measurements are frequently acquired across multiple sensing units [70]. In practical deployments, a fraction of these sensors may be faulty or compromised [64, 68], leading to a corresponding fraction of corrupted measurements. While these applications domain provide concrete instances of partial adversarial contamination, our goal is to emphasize that the proposed model (1.2) captures both gross, correlated corruptions and uncorrelated stochastic noise in a unified framework. Our corruption model is designed to abstract out more general forms of grossly correlated corruption present in real-world data.
@@ 2 Techniques
Prior techniques for phase retrieval with strong adversarial corruption and heavy-tailed noise (1.2): Naturally, the main algorithmic principle of [7] and [25] is to use a robust spectral initialization followed by a robust gradient descent procedure. We first describe their robust gradient descent procedure: they express the gradient as a mean of a random variable whose samples can be obtained from the 1.2, and leverage recent algorithmic breakthroughs for robust mean estimation; this idea has appeared in several recent works [57, 44, 39]. For robust spectral initialization, [7] have the insight that the top eigenvector of $\operatorname{Cov}(ya)$ also suffices for spectral initialization and use the exponential time robust covariance estimator [54] to estimate $\operatorname{Cov}(ya)$, and find the top eigenvector of the estimate. On the other hand [25] (who focuses on the $\sigma=0$ case) showed that given sufficient samples, robustly finding the top eigenvector of $\mathbb{E}[yaa^{T}]$ can be reduced to a convex Ky-fan-2 norm optimization problem, which can be converted into a packing semidefinite program for which nearly-linear time solvers exist. Their reduction for near-linear sample complexity has a bug, which is later claimed to be resolved in [26]. For details, see page 6, and for a full discussion, see Section 4.2.666However, their reduction works if the sample complexity is allowed to be more than $\tilde{O}(n)$ (but still a polynomial in $n$), we elaborate on this in Section 4.2.
@@ 2.1 Our technical contribution:
Comparison of our technique with prior use of truncation+filtering. While truncation and filtering are classical tools in robust statistics, their successful interplay is highly problem-dependent. Several examples illustrate this point: for instance, [20]’s robust sparse mean estimation proceeds by projecting (i.e. truncating) means to an $\ell_{\infty}$ ball around the median-of-means estimator, followed by the filtering procedure of [2]. Catoni and Olivier [11] use a customized truncation-based estimator for heavy-tailed mean estimation. The work [60] addresses the robust sparse regression problem by first using coordinate-wise truncation together with a filtering step to remove adversarially corrupted design vectors, and then they perform regression by minimizing the Huber loss, which is essentially a truncated $\ell_{2}$ loss, to handle corruption in the responses. On the other hand, the proof attempt in the (incorrect) submission [25] uses truncation together with concentration inequalities, which is conceptually similar to the truncation-plus-filtering framework, where the validity of their argument appears to rely critically on the last inequality on page 14 of their manuscript, which claims that $\operatorname{Pr}_{x\sim\mathcal{N}(0,1)}(x^{4}\geq\ln z/c)\leq z^{-3}$ holds for all $z>1$ for some constant c; however, this is false, and more generally the integral $\int^{\infty}_{\tilde{\tau}}\operatorname{Pr}_{x\sim\mathcal{N}(0,1)}(x^{4}\geq\ln z/c)dz$ diverges for all $\tilde{\tau}$ and constants $c>0$; from this viewpoint, the issues in their argument can be interpreted as a failed attempt to apply a truncation-based method, whereas in a concurrent work [26], the authors of [25] claim a result that resolves the incorrect part of their earlier results for phase retrieval in the zero-noise setting using a sophisticated truncation+filtering techniques. These examples highlight that the correct use of truncation and filtering is subtle and depends critically on the specifics of the problem: over-truncation can destroy essential structure, while under-truncation can render filtering ineffective. We truncate the vectors $ya$ using a norm-based cutoff determined by the input parameters, chosen large enough to keep the top eigenvector of the covariance of the truncated $ay$ close to the top eigenvector of the true covariance, yet small enough to apply the robust PCA method of [41]. Furthermore, unlike earlier truncation-based phase retrieval methods—which typically truncate only on the responses—our method truncates both the responses and the sensing vectors. We also show that our method remains valid even when the noise has a nonzero mean, which requires robustly solving a special case of blind deconvolution. Our work provides a principled and effective initialization scheme using truncation + filtering that may naturally extend to other robust non-convex estimation settings.
@@ 3.2 Robust principal component analysis
Note that Algorithm 2 from [41] is actually designed to return the robust top eigenvector in the case $k=1$ but we can estimate the corresponding eigenvalue also. Specifically, Algorithm 2 takes as input the corrupted set $T=\{X_{j}X_{j}^{\top}\}_{i=1}^{m}$, together with additional parameters such as the corruption level and failure probability. It then applies a filtering procedure (Algorithm 3, Double Filtering, from [41]) to remove points in $T$ that are likely to be adversarially corrupted. The average of the remaining points in $T$ is computed, which we denote by $\hat{M}$. Finally, the top eigenvector of $\hat{M}$ is returned as an estimate of the robust PCA solution, and the guarantee is given by Equation 3. In the last step, one may also compute the corresponding eigenvalue of this top eigenvector, which serves as an estimate of the leading eigenvalue of $\Sigma_{r}$. While Proposition 2.6 from [41] provides guarantees only for the eigenvector estimate (Equation 3) returned by Algorithm 2, the corresponding eigenvalue estimate can also be analyzed using the results in [41]. Its guarantee is stated in Equation 4. We provide a rigorous proof for Equation 4 in Section D.1.
@@ 4 Robust Phase Retrieval with zero mean noise
$\frac{dist(x_{P},x^{*})}{\|x^{*}\|}$ $\leq C_{5}\exp\!\left(-C_{6}P(1-\sqrt{\epsilon})\right)+C_{5}\frac{\sigma}{\|x^{*}\|^{2}}\left(\sqrt{\tfrac{n\log n}{\tilde{m}}}+\sqrt{\tfrac{\log(1/\delta)}{\tilde{m}}}+\sqrt{\epsilon}\right)$ (5)
@@ 4.1 Robust Spectral initialization
(Robust spectral initialization for phase retrieval.) Under the setting and notations given in Algorithm 1, if $\epsilon\leq C_{1}(r_{up}^{2}+1)^{-2}$, $m_{1}\geq C_{2}\log(2/\delta)(1+r_{up}^{2})$ and
$m_{2}\geq C_{3}\left(n(r_{up}^{2}+1)\log(2n/\delta\epsilon)/\epsilon\right)$. Then, with probability at least $1-\delta$, $dist\left(x_{0},x^{*}\right)\leq\left\|x^{*}\right\|/9.$ The time complexity of Algorithm 1 is $O(T_{\mathrm{rob\text{-}mean}}(m_{1},1)+m_{2}^{2}n).$
@@ 4.2 Connection between stability-based robust PCA and robust phase retrieval
By Proposition 4.7, we conclude that to obtain nearly-linear time algorithms with sample complexity $\tilde{O}(n)$ for robust phase retrieval, it suffices to prove the stability condition (Definition 4.6) for the empirical distribution with $\widetilde{O}(n)$ truncated samples, specifically, to answer in the affirmative the following question: 777[25] claims a proof of the open question for the noiseless case ($\zeta=0$) in their manuscript. However, their argument is invalid, as discussed on page 6.
@@
To answer the question partly, we show that the empirical distribution over $\tilde{O}(n^{5})$ samples of the random variable $Y$ satisfies the stability definition Definition 4.6. We prove this formally in Lemma F.3.888In a concurrent work [26], the authors of [25] claim to show that the truncated empirical distribution satisfies the stability condition (Definition 4.6) with $\tilde{O}(n)$ samples, providing an affirmative answer to 4.8 in the noise-free ($\zeta=0$) setting. Lemma F.3 leads to the Theorem F.5 (which can be found with proof in Section F.4), which shows that the algorithm $A_{alt}$ suffices for robust spectral initialization (which, along with Theorem 4.2 gives an algorithm for robust phase retrieval) and runs in $\tilde{O}(T_{\mathrm{rob\text{-}mean}}(m_{1},1)+m_{2}n)$ time but at the cost of higher sample complexity $m_{0}=m_{1}+m_{2}=\Omega(n^{5}).$
@@ 6 Conclusion
7 Acknowledgment
We thank Pravesh Kothari for useful discussions during the authors’ visit to the International Centre for Theoretical Sciences, Tata Institute of Fundamental Research (ICTS-TIFR), Bangalore, for the conference GEOMETRY, PROBABILITY, AND ALGORITHMS. We also acknowledge the support of the Department of Atomic Energy, Government of India, under Project No. RTI4014.
@@ D.1 Proof of Equation 4
As the Remark 3.3 suggests that Algorithm 2 [41] has been designed to retrieve the top eigenvector for the case $k=1$. Here , we rigorously proof Equation 4. Before proving Equation 4, we state a lemma from [41] that will be used in the proof. The lemma is as follows:
@@ Appendix E Robust Phase retrieval under Zero-Mean Noise
$\mathbb{E}\left[\left(a^{\top}x^{*}\right)^{8}\left(v^{\top}a\right)^{4}\right]\leq 10395\,\|x^{*}\|^{4}(v^{\top}x^{*})^{4}+5985\|x^{*}\|^{8}\,\|v\|^{4}\leq 16380\,\|x^{*}\|^{8}\|v\|^{4}=O(\|x^{*}\|^{8}\|v\|^{4}),$ (18)
@@
$\|\mathbb{E}[XX^{\top}\mathbb{I}_{\|X\|\geq\tau}]\|_{\mathrm{op}}$ $=\sup_{\|v\|=1}\mathbb{E}[\langle v,X\rangle^{2}\mathbb{I}_{\|X\|\geq\tau}]\overset{(a)}{\leq}\sup_{\|v\|=1}\sqrt{p}\sqrt{\mathbb{E}[\langle v,X\rangle^{4}]}$
$\leq\sqrt{p}\sup_{\|v\|=1}\sqrt{\mathbb{E}[8(\zeta^{4}+\langle a,x^{*}\rangle^{8})\langle a,v\rangle^{4}]}\overset{(b)}{\leq}\sqrt{p}\sqrt{8(3K_{4}^{4}+16380\|x^{*}\|^{8})}$
$=O\left(\sqrt{p}\sqrt{(K_{4}^{4}+\|x^{*}\|^{8})}\right)=O\left(\sqrt{p}(K_{4}^{2}+\|x^{*}\|^{4})\right),$
@@
$\operatorname{dist}(\hat{u},u)=\min\{\|\hat{u}-u\|_{2},\|\hat{u}+u\|_{2}\}\leq\sqrt{2-2\sqrt{1-O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)}},$
@@
$\geq 10\|x^{*}\|^{4}(1-\langle\hat{u},u\rangle^{2}).$
So now,
$(1-\langle\hat{u},u\rangle^{2})=O\left(\epsilon\left(1+\frac{\sigma^{2}}{\|x^{*}\|^{4}}\right)+73(r_{up}^{2}+1)\sqrt{\epsilon}\right)=O\left(\epsilon(r_{up}^{2}+1)+(r_{up}^{2}+1)\sqrt{\epsilon}\right)=O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right).$
@@
$|\langle\hat{u},u\rangle|\geq\sqrt{1-O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)}.$
Now,
$\operatorname{dist}(\hat{u},u)$ $=\min\{\|\hat{u}-u\|_{2},\|\hat{u}+u\|_{2}\}$
$=\sqrt{2-2|\langle\hat{u},u\rangle|}$
$\leq\sqrt{2-2\sqrt{1-O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)}}.$
@@
$\operatorname{dist}(\hat{u},u)=\min\{\|\hat{u}-u\|_{2},\|\hat{u}+u\|_{2}\}\leq\sqrt{2-2\sqrt{1-O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)}}\leq 1/27$
@@
$dist(\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{12\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27},\quad\text{and}$
$dist(-\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{12\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27}.$
@@
$\operatorname{dist}\left(x_{0},x^{*}\right)$ $=\operatorname{dist}\left(\hat{\|x^{*}\|}\hat{u},x^{*}\right)\leq\operatorname{dist}(\hat{\|x^{*}\|}\hat{u},\|x^{*}\|\hat{u})+\operatorname{dist}(\|x^{*}\|\hat{u},\|x^{*}\|u)+dist\left(\|x^{*}\|u,x^{*}\right)$
$=\|\hat{\|x^{*}\|}\hat{u}-\|x^{*}\|\hat{u}\|+\operatorname{dist}(\|x^{*}\|\hat{u},\|x^{*}\|u)+dist\left(\|x^{*}\|u,x^{*}\right)$
$=|\hat{\|x^{*}\|}-\|x^{*}\||+\|x^{*}\|\operatorname{dist}(\hat{u},u)+dist\left(\|x^{*}\|u,x^{*}\right)$